rationalizing the denominator. Simplify the expression. Square root of 7/8
Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

SQRT7/8=1/2SQRT7/4

Answer by stanbon(75887)

Simplify the expression. Square root of 7/8
You want the denominator to be a perfect square; so you multiply numerator
and denominator by 2 to get:
sqrt(14/16) =(sqrt14)/(sqrt16) = (sqrt14)/4
=============
Cheers,
Stan H.